OpenStudy (anonymous):

The time it takes a 800 liter tank to empty, through a pipe at the bottom of the tank, is 45 minutes. According to Toricelli's Law, the volume, V, of water remaining in the tank after t minutes is given by V(t) = 800 (1- (t/45))^2. Find V^-1 as a function of V.

OpenStudy (amistre64):

the inverse of V?

OpenStudy (amistre64):

as a function of V ?

OpenStudy (anonymous):

Yes I need to find the inverse of V as a function of V

OpenStudy (anonymous):

Yes I need to find the inverse of V as a function of V

OpenStudy (amistre64):

V(t) = 800 (1- (t/45))^2 v = 800(1-(t/45))^2 v/800 = (1-t/45)^2 sqrt(v/800) = 1- t/45 sqrt(v/800) - 1 = -t/45 45 (sqrt(v/800)-1) = -t t = -45 [sqrt(v\800)-1] thats the best I can do fer ya

OpenStudy (anonymous):

Thanks thats what I got!

OpenStudy (amistre64):

youre welcome :)